# LLM Watermark Evasion via Bias Inversion

### v3

@@ 6 Conclusion
This paper exposes fundamental vulnerabilities in LLM watermarking through a theoretical analysis, from which we developed the Bias-Inversion Rewriting Attack (BIRA). Our attack erases the watermark’s statistical signal by applying a negative logit bias to tokens identified using self-information. We empirically demonstrate that BIRA consistently evades detection from recent watermarking schemes while preserving the original text’s meaning. Our work reveals significant limitations in current methods, highlighting the need for more rigorous evaluation of watermarking and motivating the defenses that remain robust against sophisticated paraphrasing attacks.
Appendix A Proof of Theorem
Proof of Theorem 1.
@@ Appendix A Proof of Theorem
Proof of Theorem 2.
By Theorem 1, for fixed $N$ there exists $p_{\tau}$ with $\mathcal{D}(y,\mathcal{W}_{k})=\mathbf{1}\{\hat{p}(y;\mathcal{W}_{k})\geq p_{\tau}\}$. Define the indicator variables and their conditional expectations:
@@
A.1 Robustness to Proxy Green Sets
In the black-box setting, the adversary does not have access to the true green sets $\mathcal{G}(\mathcal{W}_{k})$, so we use a proxy $\widehat{\mathcal{G}}$ that may not contain all green tokens. The guarantee of Theorem 2 still holds if the following two conditions are met.
Average miss rate bound. The average probability that a sampled token is a true green token not included in the proxy is at most $\varepsilon$:
$\frac{1}{N}\sum_{n=1}^{N}\Pr\!\left(\tilde{y}^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\setminus\widehat{\mathcal{G}}\,\middle|\,\tilde{y}^{0:n-1}\right)\leq\varepsilon.$
Average proxy suppression. The attack suppresses tokens from the proxy on average, such that
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\!\left[\mathbf{1}\{\tilde{y}^{(n)}\in\widehat{\mathcal{G}}\}\,\middle|\,\tilde{y}^{0:n-1}\right]\leq p^{\prime}_{\tau}-\delta\quad\text{for some }\delta>0,$
where $p^{\prime}_{\tau}\coloneqq p_{\tau}-\varepsilon$.
Under these conditions, the conclusion of Theorem 2 still holds. This follows by showing that its premise is satisfied. Let $\bar{p}_{N}$ be the average conditional green probability:
$\bar{p}_{N}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\!\left[\mathbf{1}\{\tilde{y}^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\}\,\middle|\,\tilde{y}^{0:n-1}\right]$
$=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\!\left[\mathbf{1}\{\tilde{y}^{(n)}\in\widehat{\mathcal{G}}\}\,\middle|\,\tilde{y}^{0:n-1}\right]+\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\!\left[\mathbf{1}\{\tilde{y}^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\setminus\widehat{\mathcal{G}}\}\,\middle|\,\tilde{y}^{0:n-1}\right]$
$\leq(p^{\prime}_{\tau}-\delta)+\varepsilon$
$=(p_{\tau}-\varepsilon-\delta)+\varepsilon=p_{\tau}-\delta.$
This satisfies the premise of Theorem 2, so the result holds.
Appendix B Details of the Text Degeneration Detection Function
@@
1:Paraphrased text $\tilde{y}=[\tilde{y}^{(0)},\ldots,\tilde{y}^{(L-1)}]$; collapse window $h\in\mathbb{N}$; collapse threshold $\rho\in(0,1]$.
@@
10:  return True $\triangleright$ degeneration detected
@@
For paraphrasing, we set the maximum generation length of the LLM to 1,500 tokens. We observed that paraphrased text is typically generated normally when no degeneration occurs. However, when degeneration does occur, the text begins normally but then suddenly repeats the same phrase until the maximum token limit is reached, as shown in Appendix K.2. In the degenerated samples we examined, both the starting point of the repetition and the length of the repeated phrase varied. To ensure a sufficient detection window, we chose a large maximum generation length and a window size of $h=450$ tokens. We set the threshold $\rho=0.25$, meaning that if more than 75% of the tokens within the detection window are duplicates (which is not normal for natural text), the text is considered largely repetitive and redundant.
Figure 6: Detection upper bounds per sample from Theorem 2, sorted by BIRA. BIRA substantially reduces the upper bound on detection probability for most samples compared to Vanilla.
Appendix C Detection Bound Analysis
To validate Theorem 2, we compute the per-sample upper bound on the detection probability under the Unigram watermark for 500 samples, using the experimental setup described in Section 5.1. For ease of analysis, we generate watermarked text with Llama-3.2-3B and apply the BIRA attack with Llama-3.1-8B, since the two models share the same tokenizer. For each attacked sequence of length $N$, we calculate the average conditional green probability
set $\hat{\delta}=\max\{0,\,p_{\tau}-\bar{p}\}$, and evaluate the bound $\exp\!\big(-\tfrac{1}{2}N\hat{\delta}^{2}\big)$. Figure 6 shows that BIRA yields much lower per-sample detection bounds than Vanilla for most samples. At the 90th percentile, the upper bound for BIRA is $7.50\times 10^{-2}$, whereas for Vanilla it is $2.23\times 10^{-1}$.
@@ Appendix C Detection Bound Analysis
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP Avg ASR
Vanilla (Llama-3.1-8B) 81.0% 65.0% 70.0% 78.0% 99.0% 49.0% 68.0% 72.9%
DIPPER-1 94.0% 61.0% 73.0% 96.0% 99.0% 62.0% 87.0% 81.7%
DIPPER-2 98.0% 74.0% 83.0% 94.0% 100.0% 66.0% 95.0% 87.1%
SIRA (Llama-3.1-8B) 97.0% 91.0% 90.0% 99.0% 99.0% 73.0% 94.0% 91.9%
BIRA (Llama-3.1-8B, ours) 100.0% 99.0% 99.0% 100.0% 99.0% 98.0% 97.0% 98.9%
@@
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP Avg ASR
Vanilla (Llama-3.1-8B) 76.0% 63.0% 61.0% 83.0% 97.0% 46.0% 64.0% 70.0%
DIPPER-1 87.0% 52.0% 75.0% 92.0% 100.0% 67.0% 83.0% 79.4%
DIPPER-2 93.0% 63.0% 76.0% 95.0% 100.0% 80.0% 96.0% 86.1%
SIRA (Llama-3.1-8B) 95.0% 90.0% 90.0% 99.0% 97.0% 71.0% 88.0% 90.0%
BIRA (Llama-3.1-8B, ours) 98.0% 100.0% 96.0% 99.0% 100.0% 99.0% 95.0% 98.1%
@@
Appendix D Extended Experimental Results on Additional Datasets
To strengthen the evaluation of our method, we conduct additional experiments on two datasets, Dolly CW Conover et al. (2023) and DBPEDIA Class Raffel et al. (2020), using 100 examples and the experimental setup described in Section 5.1 with the Llama 3.1 8B model. We evaluate attack performance using the default hyperparameters, report TPR at FPR levels of 1% and 10% to mitigate the effect of a fixed threshold, and include text quality evaluation.
Figure 7 and Figure 8 show that our method consistently outperforms the baselines under the default settings. Figure 9 and Figure 10 further demonstrate that our method achieves superior performance under the adaptive threshold setup.
@@ Appendix D Extended Experimental Results on Additional Datasets
Appendix E Applicability to Sentence-Level Watermarking
Our method is primarily designed for token-level watermarking schemes, since they are fundamental and dominant paradigms. To strengthen robustness against paraphrasing, recent works Hou et al. (2024); Dabiriaghdam and Wang (2025) have proposed more robust sentence-level watermarking schemes, albeit with much higher computational cost. However, these schemes are conceptually similar to token-level watermarking, so the core principle behind BIRA naturally extends to these sentence-level methods as well.
Conceptually, sentence-level watermarking schemes Hou et al. (2024); Dabiriaghdam and Wang (2025) replace the “green/red” token distinction with a “valid/invalid” region in sentence embedding space. Let $s_{n}$ denote the $n$-th sentence and $e_{n}=f_{\text{em}}(s_{n})$ its embedding under an embedding model $f_{\text{em}}$. During generation, these methods repeatedly regenerate $s_{n+1}$ until the similarity
@@ Appendix E Applicability to Sentence-Level Watermarking
Extension of Theorem 2 to Sentence-Level Watermarking. Our Theorem 2 can be extended by redefining the random variable from an indicator of sampling a green token to an indicator of a valid sentence pair. Specifically, instead of
@@
which effectively reduces the probability that $\tilde{Z}_{n}=1$. Following the same logic as in Theorem 2, ensuring
@@
Experimental results of BIRA on sentence-level watermarking. To evaluate the effectiveness of our token-level BIRA attack against sentence-level watermarking, we conduct experiments under the setup in Section 5.1 using the Llama-3.1-8B model with the SimMark watermarking scheme Dabiriaghdam and Wang (2025), which is a state-of-the-art sentence-level watermark. For detection, we use a $z$-score threshold of $5.03$, which yields an FPR of $1\%$, and adopt the hyperparameter setting recommended by Dabiriaghdam and Wang (2025), using the interval $[0.68,0.76]$ for cosine similarity on the C4 dataset.
Table 7 shows that BIRA still achieves strong attack performance, albeit slightly weaker than on token-level watermarking schemes. We attribute this to BIRA’s strengthened paraphrasing ability, which alters token-level choices sufficiently to shift sentence embeddings outside the valid region. Interestingly, DIPPER-1 and DIPPER-2 exhibit very weak attack performance under SimMark, highlighting the need for stronger stress-testing methods for evaluation, while SIRA attains the highest ASR but degrades text quality more than BIRA. We hypothesize that SIRA’s performance gain partly arises from generating lower-quality paraphrases whose sentence embeddings deviate further from the original, whereas BIRA achieves a more favorable trade-off between watermark evasion and semantic fidelity.
@@
Appendix F Analysis of Perplexity Increase under BIRA
@@
Human evaluation of attacked text with high PPL. Qualitative analysis in Appendix K.3 shows that attacked texts with high PPL remain grammatical and reasonably readable despite their elevated perplexity scores. Our evaluation suggests that this PPL increase may stem from stylistic factors such as overuse of hyphens or particular synonym choices. Crucially, we also observe that human-written text from the C4 dataset can exhibit similarly high PPL, particularly in informal or conversational sentences (Figure 11). This supports the view that PPL is not a perfect proxy for human-perceived quality in real-world scenarios.
Can PPL be used as a post-hoc signal to flag attacked text? Perplexity alone is not a sufficiently reliable detection signal, due to its significant overlap between natural and attacked text. As illustrated in Figure 12, the PPL distributions of attacked text by BIRA for KGW watermarking and natural text in C4 dataset overlap substantially for Llama-3.1-8B and Llama-3.1-70B, making PPL ineffective for discriminating between them. For GPT-4o-mini, the BIRA distribution is more shifted toward higher PPL, but there is still considerable overlap with natural text. We suspect that part of this shift is caused by the mismatch between GPT-4o-mini’s tokenizer and the auxiliary tokenizer (Llama-3.1-3B) used to evaluate token self-information in our query-free attack setup.
@@
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP Avg ASR
Vanilla (Llama-3.1-8B) 88.8% 73.4% 73.4% 92.6% 99.8% 54.0% 80.6% 80.4%
Dipper-1 93.8% 61.2% 80.6% 92.8% 99.4% 55.6% 90.8% 82.0%
Dipper-2 97.2% 71.8% 85.4% 96.6% 99.2% 70.4% 97.2% 88.3%
ADA (Llama-3.1-8B) 98.0% 90.2% 88.4% 99.2% 100.0% 71.0% 96.6% 91.9%
SIRA (Llama-3.1-8B) 98.8% 95.0% 87.6% 99.8% 99.6% 72.8% 95.2% 92.7%
BIRA (Llama-3.1-8B, ours) 99.8% 99.4% 99.8% 100.0% 100.0% 99.6% 99.8% 99.8%
@@
Appendix G Comparison with Adaptive Attacks
@@
Appendix H Prompt for Semantic Judgment (GPT)
@@
Appendix I Paraphrasing Prompt
@@
Appendix J $z$-score comparison of attacks on different watermarking schemes
@@
Appendix K Qualitative examples
K.1 Examples of watermarked texts and attacked texts
@@
K.2 Examples of text degeneration
@@
K.3 Example of a stiff text in GPT-4o-mini
@@
Appendix L Detailed Experimental Results
L.1 Detailed experimental results for dynamic threshold
@@
L.2 Detailed experimental results of text quality evaluation

### v4

@@ 6 Conclusion
This paper formalizes a theoretical analysis of rewriting attacks and, based on this insight, introduces the Bias-Inversion Rewriting Attack (BIRA), which applies a negative bias to a proxy suppression set during rewriting. We empirically demonstrate that BIRA is a practical, query-free attack that effectively erases watermark signals while maintaining high semantic fidelity. Our findings reveal critical limitations in current methods, underscoring the need for rigorous stress-testing and defenses robust to sophisticated rewriting.
Impact Statement
This paper aims to rigorously stress-test the robustness of current LLM watermarking schemes. While numerous studies demonstrate robustness against simple post-editing operations, our work reveals that these defenses are not yet sufficiently stress-tested against sophisticated black-box evasion. We demonstrate that a simple, practical rewriting attack can effectively eliminate watermark signals, highlighting a critical gap in current evaluation standards. Additionally, we acknowledge the potential risks associated with releasing a successful evasion method, as it could be used to bypass detection of misused content. However, we believe that transparently identifying vulnerabilities is a prerequisite for scientific progress. By exposing these limitations, our work serves as a necessary form of adversarial “red-teaming” aiming to foster the development of more robust defenses and contribute to the realization of truly responsible AI systems.
Appendix A Use of LLMs
Appendix B Proof of Theorem
Proof of Lemma 4.1.
@@ Appendix B Proof of Theorem
Proof of Theorem 4.2.
By Lemma 4.1, for fixed $N$ there exists $p_{\tau}$ with $\mathcal{D}(y,\mathcal{W}_{k})=\mathbf{1}\{\hat{p}(y;\mathcal{W}_{k})\geq p_{\tau}\}$. Define the indicator variables and their conditional expectations:
@@
Appendix C Details of the Text Degeneration Detection Function
@@
1:Input: Paraphrased text $\tilde{y}=[\tilde{y}^{(0)},\ldots,\tilde{y}^{(L-1)}]$; collapse window $h\in\mathbb{N}$; collapse threshold $\rho\in(0,1]$.
@@
10:  return True (degeneration detected)
@@
For paraphrasing, we set the maximum generation length of the LLM to 1,500 tokens. We observed that rewritten text is typically generated normally when no degeneration occurs. However, when degeneration does occur, the text begins normally but then suddenly repeats the same phrase until the maximum token limit is reached, as shown in Appendix L.2. In the degenerated samples we examined, both the starting point of the repetition and the length of the repeated phrase varied. To ensure a sufficient detection window, we chose a large maximum generation length and a window size of $h=450$ tokens. We set the threshold $\rho=0.25$, meaning that if more than 75% of the tokens within the detection window are duplicates (which is not normal for natural text), the text is considered largely repetitive and redundant.
Appendix D Computational Efficiency Evaluation
Table 4: Average execution time (in seconds) for different attacks.
Attack Time
Vanilla $5.52\pm 2.82$
DIPPER $9.73\pm 2.24$
SIRA $8.57\pm 2.03$
BIRA (Ours) $7.95\pm 9.01$
To assess computational overhead, we measured the average execution time per attack over 500 samples using the KGW watermark under the setup in Section 5.1. All experiments were conducted on a single A6000 GPU, except for DIPPER built on T5-XXL (Raffel et al., 2020), which required two GPUs. As shown in Table 4, Vanilla is the most efficient baseline since it introduces no additional overhead. BIRA is the next most efficient, though it exhibits higher variance. This is caused by its adaptive bias procedure, which is designed to prevent text degeneration. This procedure was triggered in only $2.6\%$ of samples, and those rare cases had a much longer average runtime of $66.81$ seconds because repeated generation continued until the maximum length is reached. By contrast, the vast majority of samples ($97.4\%$) completed in a single iteration with an average of $6.38$ seconds, accounting for BIRA’s overall efficiency despite the variance introduced by a few outliers.
@@ Appendix D Computational Efficiency Evaluation
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP
Vanilla (Llama-3.1-8B) 81.0% 65.0% 70.0% 78.0% 99.0% 49.0% 68.0%
DIPPER-1 94.0% 61.0% 73.0% 96.0% 99.0% 62.0% 87.0%
DIPPER-2 98.0% 74.0% 83.0% 94.0% 100.0% 66.0% 95.0%
SIRA (Llama-3.1-8B) 97.0% 91.0% 90.0% 99.0% 99.0% 73.0% 94.0%
BIRA (Llama-3.1-8B, ours) 100.0% 99.0% 99.0% 100.0% 99.0% 98.0% 97.0%
@@
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP
Vanilla (Llama-3.1-8B) 76.0% 63.0% 61.0% 83.0% 97.0% 46.0% 64.0%
DIPPER-1 87.0% 52.0% 75.0% 92.0% 100.0% 67.0% 83.0%
DIPPER-2 93.0% 63.0% 76.0% 95.0% 100.0% 80.0% 96.0%
SIRA (Llama-3.1-8B) 95.0% 90.0% 90.0% 99.0% 97.0% 71.0% 88.0%
BIRA (Llama-3.1-8B, ours) 98.0% 100.0% 96.0% 99.0% 100.0% 99.0% 95.0%
@@
Appendix E Extended Experimental Results on Additional Datasets
To strengthen the evaluation of our method, we conduct additional experiments on two datasets, Dolly CW (Conover et al., 2023) and DBPEDIA Class (Raffel et al., 2020), using 100 examples and the experimental setup described in Section 5.1 with the Llama 3.1-8B model. We evaluate attack performance using the default hyperparameters, report TPR at FPR levels of 1% and 10% to mitigate the effect of a fixed threshold, and include text quality evaluation.
Table 5 and Table 6 show that our method consistently outperforms the baselines under the default settings. Figure 9 and Figure 10 further demonstrate that our method achieves superior performance under the adaptive threshold setup.
@@ Appendix E Extended Experimental Results on Additional Datasets
Appendix F Applicability to Sentence-Level Watermarking
Our method is primarily designed for token-level watermarking schemes, since they are fundamental and dominant paradigms. To strengthen robustness against paraphrasing, recent works (Hou et al., 2024; Dabiriaghdam and Wang, 2025) have proposed more robust sentence-level watermarking schemes, albeit with much higher computational cost. However, these schemes are conceptually similar to token-level watermarking, so the core principle behind BIRA naturally extends to these sentence-level methods as well.
Conceptually, sentence-level watermarking schemes (Hou et al., 2024; Dabiriaghdam and Wang, 2025) replace the “green/red” token distinction with a “valid/invalid” region in sentence embedding space. Let $s_{n}$ denote the $n$-th sentence and $e_{n}=f_{\text{em}}(s_{n})$ its embedding under an embedding model $f_{\text{em}}$. During generation, these methods repeatedly regenerate $s_{n+1}$ until the similarity
@@ Appendix F Applicability to Sentence-Level Watermarking
Extension of Theorem 4.2 to Sentence-Level Watermarking. Our Theorem 4.2 can be extended by redefining the random variable from an indicator of sampling a green token to an indicator of a valid sentence pair. Specifically, instead of
@@
which effectively reduces the probability that $\tilde{Z}_{n}=1$. Following the same logic as in Theorem 4.2, ensuring
@@
Experimental results of BIRA on sentence-level watermarking. To evaluate the effectiveness of our token-level BIRA attack against sentence-level watermarking, we conduct experiments under the setup in Section 5.1 using the Llama-3.1-8B model with the SimMark watermarking scheme (Dabiriaghdam and Wang, 2025), which is a state-of-the-art sentence-level watermark. For detection, we use a $z$-score threshold of $5.03$, which yields an FPR of $1\%$, and adopt the hyperparameter setting recommended by Dabiriaghdam and Wang (2025), using the interval $[0.68,0.76]$ for cosine similarity on the C4 dataset.
Table 7 shows that BIRA still achieves strong attack performance, albeit slightly weaker than on token-level watermarking schemes. We attribute this to BIRA’s strengthened rewriting ability, which alters token-level choices sufficiently to shift sentence embeddings outside the valid region. Interestingly, DIPPER-1 and DIPPER-2 exhibit very weak attack performance under SimMark, highlighting the need for stronger stress-testing methods for evaluation, while SIRA attains the highest ASR but degrades text quality more than BIRA. We hypothesize that SIRA’s performance gain partly arises from generating lower-quality paraphrases whose sentence embeddings deviate further from the original, whereas BIRA achieves a more favorable trade-off between watermark evasion and semantic fidelity.
@@
Appendix G Analysis of Perplexity Increase under BIRA
@@
Human evaluation of attacked text with high PPL. Qualitative analysis in Appendix L.3 shows that attacked texts with high PPL remain grammatical and reasonably readable despite their elevated perplexity scores. Our evaluation suggests that this PPL increase may stem from stylistic factors such as overuse of hyphens or particular synonym choices. Crucially, we also observe that human-written text from the C4 dataset can exhibit similarly high PPL, particularly in informal or conversational sentences (Figure 11). This supports the view that PPL is not a perfect proxy for human-perceived quality in real-world scenarios.
Can PPL be used as a post-hoc signal to flag attacked text? Perplexity alone is not a sufficiently reliable detection signal, due to its significant overlap between natural and attacked text. As illustrated in Figure 12, the PPL distributions of attacked text by BIRA for KGW watermarking and natural text in C4 dataset overlap substantially for Llama-3.1-8B and Llama-3.1-70B, making PPL ineffective for discriminating between them. For GPT-4o-mini, the BIRA distribution is more shifted toward higher PPL, but there is still considerable overlap with natural text. We suspect that part of this shift is caused by the mismatch between GPT-4o-mini’s tokenizer and the auxiliary tokenizer (Llama-3.1-3B) used to evaluate token self-information, and GPT-4o-mini’s writing property that generate lots of hyppen.
@@
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP
Vanilla (Llama-3.1-8B) 88.8% 73.4% 73.4% 92.6% 99.8% 54.0% 80.6%
Dipper-1 93.8% 61.2% 80.6% 92.8% 99.4% 55.6% 90.8%
Dipper-2 97.2% 71.8% 85.4% 96.6% 99.2% 70.4% 97.2%
ADA (Llama-3.1-8B) 98.0% 90.2% 88.4% 99.2% 100.0% 71.0% 96.6%
SIRA (Llama-3.1-8B) 98.8% 95.0% 87.6% 99.8% 99.6% 72.8% 95.2%
BIRA (Llama-3.1-8B, ours) 99.8% 99.4% 99.8% 100.0% 100.0% 99.6% 99.8%
@@
Appendix H Comparison with Adaptive Attacks
@@
Appendix I Prompt for Semantic Judgment (GPT)
@@
Appendix J Rewriting Prompt
@@
Appendix K $z$-score comparison of attacks on different watermarking schemes
@@
Appendix L Qualitative examples
L.1 Examples of watermarked texts and attacked texts
@@
L.2 Examples of text degeneration
@@
L.3 Example of a stiff text in GPT-4o-mini
@@
Appendix M Detailed Experimental Results
M.1 Detailed experimental results for dynamic threshold
@@
M.2 Detailed experimental results of text quality evaluation